OpenStudy (anonymous):

Simplify: (9x^3y^4-3x^2y^3) / 3x^2y^3 A. 3xy - 1 B. 9x^3y^4 C. 6xy D.2x^2y^3 My answer: A but I need someone to show me how to get the answer my teacher not showing me how to get it.

OpenStudy (anonymous):

@Rubyjadeluv

OpenStudy (kendricklamar2014):

You are correct. The answer is A.

OpenStudy (kendricklamar2014):

Here are the steps to solve this problem: \[\frac{ 9x^3y^4−3x^2y^3 }{ 3x^2y^3 } = \frac{ 9xy−3 }{ 3 }\]

OpenStudy (kendricklamar2014):

Now, we simplify again:\[\frac{ 9xy−3 }{ 3 } = 3xy-1\]

OpenStudy (kendricklamar2014):

Hope it helped @Takirap2000

OpenStudy (anonymous):

thanks so much kendrick

OpenStudy (kendricklamar2014):

You're Welcome